Obtain or compute the following quantities.

(a)

F0.05, 4, 7

(Round your answer to two decimal places.)

(b)

F0.05, 7, 4

(Round your answer to two decimal places.)

(c)

F0.95, 4, 7

(Round your answer to three decimal places.)

(d)

F0.95, 7, 4

(Round your answer to three decimal places.)

(e)

the 99th percentile of the F distribution with

v1 = 8,

v2 = 15

(Round your answer to two decimal places.)

(f)

the 1st percentile of the F distribution with

v1 = 8,

v2 = 15

(Round your answer to three decimal places.)

(g)

P(F ≤ 19.30)

for

v1 = 5,

v2 = 2

(Round your answer to two decimal places.)

(h)

P(0.207 ≤ F ≤ 4.00)

for

v1 = 12,

v2 = 6

(Round your answer to two decimal places.)

You may need to use the appropriate table in the Appendix of Tables to answer this question.
Answer #1

Solution

We will find the solution using R

> a=qf(0.05,4,7)
> round(a,2)
[1] 0.16
> b=qf(0.05,7,4)
> round(b,2)
[1] 0.24
> c=qf(0.95,4,7)
> round(c,2)
[1] 4.12
> round(c,3)
[1] 4.12
> c
[1] 4.120312
> d=qf(0.95,7,4)
> round(d,3)
[1] 6.094
> g=pf(19.30,5,2)
> round(g,2)
[1] 0.95
> h=pf(4,12,6)-pf(0.207,12,6)
> round(h,2)
[1] 0.94